OpenStudy (vishweshshrimali5):

Integrate cos x cos 2x cos 3x w.r.t x

OpenStudy (anonymous):

um guess u shud do it the long way

OpenStudy (vishweshshrimali5):

which one ?

OpenStudy (vishweshshrimali5):

u means the first expanding cos 2x and cos 3x and then using the product rule

OpenStudy (anonymous):

wait cos2x= 2sin^2x -1

OpenStudy (anonymous):

cos3x = cos(2x+x) expand it

OpenStudy (anonymous):

nd put sinx=t

OpenStudy (vishweshshrimali5):

ok

sam (.sam.):

Use \[\cos (A ) \cos (B )=\frac{1}{2} (\cos (A -B )+\cos (A+B))\]

sam (.sam.):

\[\frac{1}{2}\int\limits \cos ^2(3 x) \, dx+\frac{1}{4}\int\limits (\cos (2 x)+\cos (4 x)) \, dx\]

OpenStudy (anonymous):

i missed that :P

OpenStudy (vishweshshrimali5):

Ok got that. but what about cos (2x) and cos (4x)

sam (.sam.):

u=2x h=4x t=3x du=2dx dh=4dx dt=3dx \[\frac{1}{8}\int\limits \cos (u) \, du+\frac{1}{6}\int\limits \cos ^2(t) \, dt+\frac{1}{16}\int\limits \cos (h) \, dh\]

OpenStudy (vishweshshrimali5):

ok that would work for sure. Thanks a lot for help

sam (.sam.):

welcome :)